### v1

# Article Type: Pre-Print

Abstract
Cybersecurity threats in automotive systems pose significant risks to safety and reliability. This article introduces a methodology integrating threat-informed dynamic security modelling with a Threat Analysis and Risk Assessment workflow. Using the example of an In-Vehicle Infotainment system, we demonstrate the methodology’s application in risk management to strengthen automotive resiliency.
Keywords: risk management, security, autonomous vehicles
journal: IEEE Security & Privacy
Introduction
Automotive systems have transformed into sophisticated platforms brimming with cutting-edge features. In contrast to their predecessors, today’s automotive systems are equipped with a variety of advanced sensors and connectivity options, such as Control Area Networks (CAN), Local Interconnect Networks (LIN), FlexRay, and Ethernet. They also offer seamless external connectivity through Bluetooth, Wi-Fi, and GPS. Additionally, the integration of feature-rich In-Vehicle Infotainment (IVI) systems and powerful computing units for autonomous driving continues to push the boundaries of what is possible in automotive technology. However, this technological evolution also broadens the attack surface, making automotive systems more susceptible to cybersecurity attacks.
Cybersecurity threats and vulnerabilities in the automotive industry are well-documented, and attackers exploit these vulnerabilities to compromise vehicle security, access sensitive data, and even take control of vehicle functions.
Recent literature has increasingly highlighted the challenges of defending against cybersecurity threats in automotive systems. In their survey on security attacks and defence techniques for Connected and Autonomous Vehicles (CAVs), Pham and Xiong (2021) observed that attackers could potentially compromise one vehicle component after another. To mitigate this risk, they emphasised the need for manufacturers to strengthen the security of all system components to ensure the overall protection of CAVs.
While this approach may be feasible during the early stages of an automotive system’s life cycle, evolving cybersecurity threats are likely to render it less effective over time. Given the rapid increase in both threats and the complexity of automotive systems, it is unrealistic to expect that an automotive system’s initial cybersecurity features will counter all emerging risks throughout its entire life cycle. This is where threat-informed dynamic threat and attack models become crucial, offering a proactive and adaptive approach to designing secure automotive systems.
Dynamic threat and attack models provide engineers with a structured framework to analyse and mitigate cybersecurity risks, ensuring that automotive systems remain protected against evolving cybersecurity threats. Additionally, they enable consumers to make informed decisions when selecting vehicles based on their cybersecurity robustness.
However, many of the threat and attack models discussed in the literature do not fully address the unique cybersecurity requirements of modern automotive systems. In our previous work [1], we conducted a systematic literature review of threat and attack modelling approaches pertinent to cyber-physical systems (CPSs)—systems that integrate computational and physical process—of which automotive systems are an example. Our investigation yielded several findings, including:
  • Threat modelling is commonly conducted only in early stages of system development, and does not always consider attack models.
  • Threat and attack models in the literature primarily focus on IT systems, making them challenging to apply to automotive systems.
  • Much of the existing literature does not differentiate between cybersecurity breaches and consequences in IT systems versus those in automotive systems.
  • Ambiguity persists in the literature regarding the definitions and relationships between threat modelling and attack modelling.
These findings highlight the need to tailor security modeling approaches for automotive systems so that they can adequately capture the unique cybersecurity threats faced in these environments.
For instance, threat models developed using Microsoft’s STRIDE (Spoofing, Tampering, Repudiation, Information Disclosure, Denial of Service, and Elevation of Privilege) primarily focus on data flows and information security, overlooking critical aspects of how such breaches can directly affect automotive functions and safety. In other words, threat modelling of automotive systems often do not consider consequences of cybersecurity attacks beyond information security. On a similar note, attack models designed with traditional IT systems in mind do not fully align with multi-path tactics and techniques employed by attackers seeking to gain control of a vehicle. These techniques often involve a combination of physical and cyber actions.
In short, our investigation has highlighted limitations in current threat and attack modelling methods that underscore the need for a tailored, dynamic approach. Threat and attack models need to iteratively and continuously consider both consequences and characteristics of cyber-physical attacks so as to offer a comprehensive approach to safeguarding automotive systems.
Contributions. In this article, we introduce ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling) — a novel security modeling approach that enhances the resilience of automotive systems by dynamically updating their cybersecurity posture in response to prevailing threats, attacker tactics, and their impact on system functionality and safety. We demonstrate the effectiveness of ACTISM by applying it to an IVI system, illustrating how the model can be adapted as new threats emerge.
Unlike existing security modelling approaches, ACTISM dynamically updates security models with new information acquired through threat intelligence and ongoing security monitoring. This keeps security models relevant against evolving attacker tactics, techniques and procedures (TTPs) even if the models were developed in early stages of system development. ACTISM also allows practitioners to apply the security models in both IT and automotive domains by using attack paths to capture the multi-path and multi-agent aspects of real-world cyber-physical attacks.
Cybersecurity incidents in automotive systems differ significantly from those in traditional IT systems. These incidents can lead to complex failure modes with both cyber and physical consequences, such as vehicle misnavigation and potentially dangerous outcomes. By adopting a consequence-driven and cyber-informed approach to cybersecurity, ACTISM ensures that cyber-physical attacks, effects and consequences are considered in security modelling for automotive systems.
ACTISM seamlessly and systematically integrates with, and adapts to a number of threat analysis and risk assessment frameworks such as:
  • Microsoft’s STRIDE methodology.
  • HEAVENS (HEAling Vulnerabilities to ENhance Software Security and Safety), a structured and model-based hazard analysis and risk assessment method for automotive systems.
  • Attack Tree, a hierarchical diagram representing potential attack paths, illustrating how an attacker could exploit system vulnerabilities to achieve specific goals.
  • Common Vulnerability Scoring System (CVSS), an open industry standard for assessing the severity of computer system security vulnerabilities.
  • Threat Analysis and Risk Assessment (TARA) methodology from the ISO/SAE 21434:2021 ("Road Vehicles - Cybersecurity Engineering") Standard, jointly developed by the International Organisation for Standardisation (ISO) and the Society of Automotive Engineers (SAE), for managing cybersecurity threats and vulnerabilities throughout the life cycle of a vehicle.
To demonstrate the practical application of ACTISM, we developed a comprehensive security model for an IVI system. We outline the step-by-step process of building the security model, emphasising the integration of threat intelligence and contextual information to accurately characterise and assess cybersecurity risks. Furthermore, we illustrate ACTISM’s adaptive response to a real-world vulnerability disclosure affecting the IVI system, underscoring its effectiveness in addressing emerging threats and showcasing its potential applicability to other automotive systems.
Finally, we outline avenues for future research and development in this area.
Related Work
Various articles have attempted to address security modelling in automotive systems.
Das et al. [2] employed the STRIDE methodology for threat modelling and utilised the SAHARA (Safety-Aware Hazard Analysis and Risk Assessment) and DREAD (Damage, Reproducibility, Exploitability, Affected Users, and Discoverability) methodologies to assess risks associated with Infotainment High Performance Computing (HPC) systems. However, their analysis did not incorporate attack modelling.
Abuabed et al. [3] proposed a framework for assessing the cybersecurity vulnerabilities of automotive vehicles. Their approach employed STRIDE to classify threats. For assessing the impact of threats, they used version 1.0 of the HEAVENS security model [4], and they evaluated attack feasibility with CVSS version 2.0. Additionally, they used a risk matrix to assess risks based on a use-case involving the Advanced Driver Assistance System (ADAS). Using a TARA approach compatible with the ISO/SAE 21434:2021 standard, the framework utilises attack trees to determine attack paths, facilitating a focused analysis of relevant threat vectors and potential attacks during the assessment of attack feasibility.
The framework has certain limitations. Firstly, the threat identification process did not include physical attack vectors, such as accessing and modifying embedded software in vehicle components, in the STRIDE Data Flow Diagram (DFD). This may arise from the assumption that physical attacks do not involve data flows. However, physical attack vectors, such as "USB Man-in-the-Middle (USB MITM)" attacks, can exploit and facilitate data flows by intercepting and altering communication. As a result, neglecting physical attack vectors introduces a significant gap in the overall security design of the vehicle.
Secondly, both the HEAVENS version 1.0 and CVSS version 2.0 frameworks adopted in Abuabed et al. [3] are no longer current. Their newer versions include improvements that deliver more granularity and accuracy in the resulting risk assessment. Lastly, the framework does not adequately address threats in later stages of the life cycle. As automotive system life cycles are typically 10 to 20 years or longer, the vehicle system could become increasingly vulnerable to new types of attacks that were not anticipated during the initial security assessment.
Chlup et al. [5] proposed an automated threat modelling tool for the automotive domain. It combines threat modelling and attack tree analysis using automatically-generated attack trees and attack paths. Additionally, it uses anti-patterns to perform in-depth analysis in the threat model. However, the approach does not perform attack feasibility estimation and risk assessment, which are key steps in the TARA workflow.
Other literature relevant to automotive cybersecurity includes a study by Girdhar et al. [6], who used the STRIDE methodology to create a threat model focused on post-accident cyberattack event analysis for CAV.
Additionally, Dang et al. [7] used a STRIDE framework to analyse and classify cybersecurity threats in an autonomous vehicle architecture, including the infotainment system. While the authors ranked the threats using CVSS version 3.0 scores, they did not assess the corresponding impact. The authors focused only on cybersecurity threats originating from outside the vehicle. This means that attacks that involve access to the vehicle’s internal architecture from within the vehicle are excluded in the analysis.
The current landscape of security modelling highlights several limitations. Firstly, there are differing assumptions regarding which vehicle assets fall within the trust boundaries of security models. Secondly, the feasibility and impact of cyber-physical attacks are rarely considered in security modelling. These gaps underscore the need for an iterative and cohesive framework to effectively address the evolving challenges in automotive cybersecurity.
Threat-informed Dynamic Security Modelling with ACTISM
We propose ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling) - a methodology for modelling security threats and attack vectors in a dynamic, iterative manner, specifically designed for automotive systems. We will illustrate the application of ACTISM to an IVI system as a running example.
Unlike conventional security modelling methodologies, ACTISM seamlessly integrates both threat and attack modelling in a continuous, iterative workflow throughout all stages of the vehicle’s life cycle. Moreover, ACTISM is dynamic, receiving updates through cybersecurity threat information acquired from vulnerability advisories and ongoing security monitoring.
Figure 1: Overview of threat-informed dynamic security modelling
As illustrated in Figure 1, a threat-informed dynamic security modeling approach relies on the seamless integration of threat and attack modeling, utilising methodologies and tools like Microsoft’s STRIDE and CVSS 3.1. A key component of this approach is the strategic incorporation of threat intelligence through continuous monitoring, which ensures that the security model remains adaptive and effective in mitigating emerging risks.
Such an iterative modelling approach is vital because threat and attack models created in the early stages of system life cycles can become outdated or irrelevant over time as new vulnerabilities emerge. This gap widens, particularly when cyber-physical attacks occur like those targeting keyless entry systems [8], breaching the trust boundary within security models.
ACTISM does not replace threat and attack modelling. Instead, it streamlines their adoption and utilisation by cybersecurity practitioners. By incorporating an iterative modelling approach, the methodology enhances the practicality and efficacy of addressing evolving cybersecurity threats and attacker capabilities. ACTISM enhances the TARA framework by dynamically updating the cybersecurity posture of automotive systems based on the evolving threat landscape and the specific impacts of potential attacks on system functionality and safety. ACTISM comprises eight key steps, outlined as follows:
  1. 1.
    Identify assets
  2. 2.
    Construct Data Flow Diagrams
  3. 3.
    Model threats
  4. 4.
    Evaluate impact
  5. 5.
    Model attacks
  6. 6.
    Evaluate attack feasibility
  7. 7.
    Evaluate risks
  8. 8.
    Review, update, and iterate
The details of each step will be elaborated on in the next section, using the running example of an IVI system.
Lastly, we further incorporate the following enhancements in ACTISM to increase its practical utility amongst cybersecurity practitioners:
  • Threat analysis: We employ the STRIDE methodology using a STRIDE-per-element approach where threats are assessed against each element in the Data Flow Diagram. We do not use a STRIDE-per-interaction approach, as this yields a statistically significant lower level of completeness compared to the STRIDE-per-element approach [9]. Additionally, we use HEAVENS version 2.0 [10] for impact assessment, as version 2.0 offers enhancements over version 1.0, providing more accurate and comprehensive risk assessments.
  • Attack analysis: We use CVSS version 3.1 to assess attack feasibility ratings, allowing us to leverage current cybersecurity vulnerability knowledge bases. Additionally, CVSS version 3.1 offers granular metrics and scoring options, which consider factors affecting the exploitability and impact of a vulnerability.
Applying ACTISM to an In-Vehicle Infotainment (IVI) System
To illustrate the practical application of ACTISM, it is essential to understand the IVI system and its role in automotive cybersecurity.
Legacy head units offered basic features such as radio reception and CD playback, with limited connectivity. In contrast, modern IVI systems are advanced digital platforms that integrate various functionalities beyond audio and navigation. These systems serve as central hubs, providing seamless connectivity to external networks, managing complex data streams, enabling access to vehicle diagnostics, and supporting interactive user interfaces for both drivers and passengers.
While modern IVI systems enhance convenience and operational efficiency, they also expose vehicles to increased cybersecurity risks. These systems are vulnerable to traditional threats, such as remote exploits and data breaches, as well as emerging threats like cyber-physical attacks targeting vehicle functionalities.
The complexity of modern IVI systems significantly expands their attack surface. They integrate multiple wireless communication protocols—Wi-Fi, Bluetooth, and cellular networks—and connect to external services and mobile devices, increasing potential vulnerabilities. Malicious actors can exploit weaknesses in wireless protocols or insecure network connections to gain unauthorised access. Furthermore, advanced functionalities like voice recognition and artificial intelligence add layers of complexity, creating additional attack vectors.
A dynamic approach like ACTISM allows for iterative security modelling throughout the IVI system’s life cycle, ensuring comprehensive protection against evolving threats.
Step 1: Identify assets
In the first step, identifying vehicle assets within the IVI system threat model typically involves a collaborative effort among system architects, engineers, and cybersecurity experts. This process integrates various activities, including reviewing technical specifications, examining system architecture diagrams, and analysing functional requirements documents. Additionally, considering the interconnected nature of modern IVI systems, potential dependencies and interactions between assets are also analysed to understand how vulnerabilities in one component could impact overall system security. These efforts aim to ensure a comprehensive understanding and coverage of all relevant assets within the IVI system.
In this article, we reference the IVI HPC system DFD from Das et al. (2024) [2] to guide our identification of IVI system assets. This initial asset identification, listed below, serves as a foundational step for conducting further detailed analysis and security modelling.
  • In-vehicle modules
    • In-Vehicle Infotainment (IVI) System
    • Human-machine Interface (HMI)
    • Central Gateway
    • GPS Receiver
    • USB
    • Bluetooth
    • Wi-Fi
  • Data stores
    • Local maps data store
    • IVI system embedded data store
Step 2: Construct Data Flow Diagrams
In the next step, we construct and analyse DFDs to identify the vehicle trust boundary and potential threat and attack vectors. This process uses the Microsoft Threat Modelling Tool (Version 7.3.31026.3) along with the Automotive Threat Modelling Template [11] developed by the NCC Group. We conduct DFD analysis for both a typical automotive system [12] and specifically for the IVI system (Figure 2) to ensure thorough identification and analysis of potential threats and vulnerabilities.
Figure 2: Data Flow Diagram for an In-Vehicle Infotainment (IVI) System
Step 3: Model threats
Next, we use the Microsoft Threat Modelling Tool to generate threats based on the previously constructed Data Flow Diagram (DFD) for the IVI system. The tool categorises these threats according to the STRIDE threat model automatically, providing an initial classification into categories such as Spoofing, Tampering, Repudiation, Information Disclosure, Denial of Service, and Elevation of Privilege.
While the tool categorises threats by default and attempts to provide threat descriptions based on its own threat catalogues and templates, the initial output may be somewhat generic and not specifically tailored to the nuances of the IVI system’s operating environment. Therefore, our analysis involves further contextualisation and systematic review of these categorised threats.
We focus on the in-vehicle modules and data stores identified in the earlier step, ensuring that the generated threats are relevant to these components. After refining the list, we update threat descriptions to better align with the specific interfaces and nuances of the IVI system. A domain expert can perform this refinement relatively easily using the Microsoft Threat Modelling Tool, taking about 30 minutes for this example. The final curated list of threats for the IVI system is shown in Figure 3.
Figure 3: Threats generated by Microsoft Threat Modelling Tool
Step 4: Evaluate impact
To analyse the impact of the identified threats, we use the HEAVENS version 2.0 framework [10] which is specifically designed for automotive risk assessments. This framework allows us to systematically evaluate the potential impacts across four key dimensions: safety (s), financial (f), operational (o), and privacy and legislation (pl).
Our approach to impact analysis is as follows:
  • Assess impact values: We assign numerical values to each threat based on its impact level on safety, financial stability, operational continuity, and compliance with privacy and legislation. (Table 1)
    Table 1: Mapping of impact levels to numerical values
    Impact Level Value
    None 0
    Low 1
    Medium 10
    High 100
  • Calculate impact score: Equation 1 is used to compute the overall impact score of each threat. This equation combines the weighted values of safety, financial, operational, and privacy and legislation impacts. To reflect the critical importance of safety in the context of automotive cybersecurity, we set $w_{s}=10$, placing the maximum weight on safety.
    Impact $\small\text{Impact Score}=\frac{w_{s}i_{s}+w_{f}i_{f}+w_{o}i_{o}+w_{pl}i_{pl}}{w_{s}+w_{f}+w_{o}+w_{pl}}$ (1)
    where $w_{s}=10$ and $w_{f}=w_{o}=w_{pl}=1$.
  • Derive impact rating: Based on the computed impact score of each threat, we derive the corresponding ratings using the thresholds defined in Table 2. This rating helps prioritise threats based on their potential severity and guides subsequent risk mitigation in our analysis.
    Table 2: Mapping of computed impact scores to ratings
    Impact Score Rating
    0.00 None
    $0.010.00<x<$ 0.01 Low
    $0.050.01\leq x<$ 0.05 Medium
    $0.450.05\leq x<$ 0.45 High
    $1.000.45\leq x<$ 1.00 Critical
The final impact evaluation is tabulated in Figure 4.
Figure 4: Impact assessment based on HEAVENS 2.0
Step 5: Model attacks
In this step, we use an attack tree to model attacker goals and map out attack paths associated with threat and attack vectors identified earlier using the Microsoft Threat Modelling Tool.
Since the vehicle driver is within the car trust boundary (Figure 2), we will assume the attacker personas and their goals as follows:
  • Persona A: A car thief interested in stealing sensitive data from the In-Vehicle Infotainment system.
  • Persona B: A rogue security researcher interested in manipulating vehicle functions that compromise the safety of the vehicle.
The attack tree depicted in Figure 5 illustrates the potential attack paths these two personas could employ to achieve their respective goals.
Figure 5: Attack tree for IVI system
Step 6: Evaluate attack feasibility
Next, we use CVSS version 3.1 to analyse and compute the attack feasibility ratings. CVSS provides a standardised method for assessing the severity of vulnerabilities based on their impact and exploitability factors.
The CVSS 3.1 base score is computed using Equation 2.
Base Score=Round up(0.6×Impact+0.4×Exploitability1.5Impact $\tiny\text{Base Score}=\text{Round up}\left(\frac{{0.6\times\text{Impact}+0.4\times\text{Exploitability}-1.5}}{{\text{Impact Factor}+0.8}}\right)$ (2)
where Impact, Exploitability, and Impact Factor are derived from the following components:
  • Impact Metrics (C, I, A):
    • Confidentiality (C): The potential impact on confidentiality.
    • Integrity (I): The potential impact on integrity.
    • Availability (A): The potential impact on availability.
  • Exploitability Metrics (AV, AC, PR, UI):
    • Attack Vector (AV): How the vulnerability is accessed.
    • Attack Complexity (AC): The complexity of the attack required to exploit the vulnerability.
    • Privileges Required (PR): The level of privileges required to exploit the vulnerability.
    • User Interaction (UI): The level of user interaction required to exploit the vulnerability.
  • Impact Factor:
    • The Impact Factor depends on the scope of the vulnerability, which can be either Scope Unchanged (U) or Scope Changed (S). If the vulnerability affects other system components, the Scope is considered Changed (S), otherwise, it is Unchanged (U).
Additionally, the CVSS temporal score takes into account factors such as exploit code maturity, remediation level, and report confidence. It is calculated using Equation 3.
Temporal Score=Round up(Base Score×Exploit Code Maturity×Remediation Level×Report $\tiny\begin{split}\text{Temporal Score}=\text{Round up}(\text{Base Score}\times\text{Exploit Code Maturity}\times\\ \text{Remediation Level}\times\text{Report Confidence})\end{split}$ (3)
Figure 6: Attack feasibility ratings based on CVSS version 3.1
Figure 6 lists the attack feasibility ratings computed using the CVSS version 3.1 calculator (https://nvd.nist.gov/vuln-metrics/cvss/v3-calculator). These ratings are based on the information available at the time of assessment during the early design stage of the IVI system.
Figure 7: Risk ratings evaluated using a 5x5 risk matrix
Step 7: Evaluate risks
In this step, we evaluate the cybersecurity risks of the IVI system using a 5x5 risk matrix. The corresponding risk ratings for identified threats are tabulated in Figure 7.
These risk ratings serve as a guideline for the Original Equipment Manufacturer (OEM) of the vehicle or the manufacturer of the IVI system. They can use these ratings to prioritise the implementation of mitigation controls during the design phase of the system life cycle. This approach ensures that adequate attention and resources are allocated to address the most critical cybersecurity concerns identified through this comprehensive analysis.
Step 8: Review, update, and iterate
In this final step, we continuously review the cybersecurity analysis and update the risk assessment based on new threat information, insights or changes in the automotive system environment. The updated risk assessment helps refine existing mitigation strategies and enhances security controls against emerging cybersecurity threats. This iterative cycle of review and improvement is crucial for maintaining the automotive system’s long-term cybersecurity posture.
Updating the Security Model After a Vulnerability Disclosure
In the earlier section, we discussed the TARA framework as part of our approach to security modelling for automotive systems. Generally, this is performed during the early stages of system life cycles. However, evolving cybersecurity threats necessitate the incorporation of new knowledge gained from ongoing security monitoring and threat intelligence throughout the system’s life cycle.
ACTISM can be seamlessly integrated with the TARA framework to create a threat-informed dynamic security modelling methodology.
We now apply this methodology to a real-world example using the Tesla Electric Vehicle (EV) jail-break vulnerability disclosure [13].
After gathering technical details from the vulnerability disclosure, we review the IVI system security model developed in the earlier section and update the CVSS metrics as necessary.
For this vulnerability disclosure, there is no need to include additional assets in the security model or to revise the DFD (Figure 2) and attack tree (Figure 5). Although there is an opportunity to enhance the DFD and attack tree to include new attack vectors, such as cyber-physical attacks like keyless entry systems [8], we will not do so in this example. However, it is critical to reassess the CVSS metrics associated with Threat ID 9146, which correlates with the vulnerability disclosure.
The CVSS metrics corresponding to Threat ID 9146 are updated as follows:
  • Scope (S) From Unchanged (UC) to Changed (C), as successful exploitation of the vulnerability could grant an attacker full control over the IVI system, potentially enabling them to access and manipulate certain vehicle functions.
  • Integrity (I) From Low (L) to High (H), as successful exploitation of the vulnerability could grant an attacker root-level read and write access to the IVI file system.
  • Availability (A) From None (N) to High (H), as an attacker could deny access to resources in the IVI system.
  • Exploit Code Maturity (E) From Unproven (U) to Proof-of-Concept (P), in view of the disclosure from the security researchers.
  • Remediation Level (RL) From Unavailable (U) to Temporary Fix (T), as the AMD Trusted Platform Module (TPM) in affected IVI systems could still be susceptible to voltage glitching attacks.
  • Report Confidence (RC) From Unknown (U) to Confirmed (C), in view of the disclosure from the security researchers.
The updated risk rating based on these revised metrics is shown in Figure 8. In summary, the corresponding risk changed from Medium to High.
This ability to perform timely updates to the risk assessment is crucial for maintaining robust cybersecurity in automotive systems. By continuously adapting to new threats and changes in attack feasibility and impact, proactive measures can be implemented to safeguard automotive system safety and functionality. This dynamic approach not only enhances the resilience of automotive systems against evolving cybersecurity threats but also reinforces the importance of informed decision-making in the face of emerging risks. Ultimately, it empowers stakeholders to take appropriate actions that protect both the automotive system and its users.
Figure 8: Updated risk rating following vulnerability disclosure
Discussion
Unlike early-stage security modelling, where OEMs and IVI system manufacturers implement comprehensive measures, addressing cyber-physical vulnerabilities later in the life cycle often involves temporary fixes or workarounds. These may not fully mitigate the risks, so vehicle owners can use risk ratings to make informed decisions. By adopting additional precautions, they can enhance the cybersecurity resilience of their vehicles throughout their life cycle.
This underscores the need for a continuous, iterative, and threat-informed security modelling process. Ongoing threat intelligence gathering, including real-time monitoring and tracking vulnerability disclosures, is crucial. A comprehensive repository of vulnerability attributes, stored in a standardised format such as JSON, enables automation. Tools can parse this data, extract relevant information, and enrich the security model via workflows like TARA.
CVSS scores and risk ratings could then be dynamically recalculated based on updated threat context. Security models can be formalised using Domain-Specific Languages (DSLs) such as the Meta Attack Language (MAL), allowing clearer descriptions of threats and vulnerabilities. Furthermore, adversary tactics from the MITRE ATT&CK knowledge base could be incorporated to enhance these models.
Machine learning approaches could also help prioritise vulnerabilities by analysing their impact, using historical data and threat intelligence to improve automotive cybersecurity.
Conclusion
Automotive systems are increasingly facing cybersecurity threats, with attackers targeting vulnerabilities that pose risks to data, functionality, and safety. Ensuring the security of automotive systems is critical to maintaining their integrity and reliability.
We introduced ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling), a novel approach designed to enhance the resilience of automotive systems by dynamically updating their cybersecurity posture. ACTISM responds to evolving threats and attacker tactics, ensuring security models remain relevant throughout a system’s lifecycle.
Unlike traditional models, ACTISM continuously integrates new data from threat intelligence and security monitoring, keeping protection up to date. It also captures the multi-agent nature of cyber-physical attacks, making it applicable to automotive domains. This integration strengthens the accuracy and completeness of security models, reflecting the complexity of real-world cyber-physical attacks.
Cybersecurity incidents in automotive systems differ from traditional IT breaches due to their potential physical consequences, such as vehicle misnavigation or compromised safety. ACTISM’s consequence-driven approach ensures that cyber-physical risks are considered, directly addressing the operational and safety integrity of automotive systems.
To demonstrate ACTISM’s practical application, we developed a comprehensive security model for an IVI system. This involved identifying key assets, conducting threat modelling, and performing risk assessments, resulting in a robust security framework. A real-world case study further highlighted ACTISM’s ability to adapt to emerging threats and its potential for broader use in other automotive technologies.
In conclusion, ACTISM offers a proactive, adaptable approach to automotive cybersecurity, addressing current challenges while preparing for the evolving landscape of automotive technology.