### v2

# ACTISM: Threat-Informed Dynamic Security Modelling for Automotive Systems

Abstract
Cybersecurity threats in complex cyber-physical systems pose significant risks to system functionality and safety. This experience report introduces ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling) — an integrated security modeling approach that enhances the resilience of automotive systems by dynamically updating their cybersecurity posture in response to prevailing threats, attacker tactics, and their impact on system functionality and safety. ACTISM addresses the existing knowledge gap in static security assessment methodologies by providing a dynamic and iterative framework. We demonstrate the effectiveness of ACTISM by applying it to a real-world example of the Tesla Electric Vehicle’s In-Vehicle Infotainment system, illustrating how the security model can be adapted as new threats emerge. We also outline avenues for future research and development in this area, including automated vulnerability management workflows for automotive systems.
Keywords:
Risk management Security Automotive systems
1 Introduction
Automotive systems have transformed into complex cyber-physical systems brimming with cutting-edge features. In contrast to their predecessors, today’s automotive systems are equipped with a variety of advanced sensors and connectivity options, such as Control Area Networks (CAN), Local Interconnect Networks (LIN), FlexRay, and Ethernet. They also offer external connectivity through Bluetooth, Wi-Fi, and GPS. In addition, the integration of feature-rich In-Vehicle Infotainment (IVI) systems and powerful computing units for autonomous driving continues to push the boundaries of what is possible in automotive technology. However, this technological evolution also broadens the attack surface, making automotive systems more susceptible to cybersecurity attacks.
Cybersecurity threats and vulnerabilities in the automotive industry are well documented, and attackers exploit these vulnerabilities to compromise vehicle security, access sensitive data, and even take control of vehicle functions.
Recent literature has increasingly highlighted the challenges of defending against cybersecurity threats in automotive systems. In their survey on security attacks and defence techniques for Connected and Autonomous Vehicles (CAVs), Pham and Xiong [11] observed that attackers could potentially compromise one vehicle component after another. To mitigate this risk, they emphasised the need for manufacturers to strengthen the security of all system components to ensure the overall protection of CAVs.
While this approach may be feasible during the early stages of an automotive system’s life cycle, evolving cybersecurity threats are likely to render it less effective over time. Given the rapid increase in both threats and the complexity of automotive systems, it is unrealistic to expect that an automotive system’s initial cybersecurity features will counter all emerging risks throughout its entire life cycle. This is where threat-informed dynamic threat and attack models become crucial, offering a proactive and adaptive approach to designing secure automotive systems.
Dynamic threat and attack models provide engineers with a structured framework to analyse and mitigate cybersecurity risks, ensuring that automotive systems remain protected against evolving cybersecurity threats. Additionally, they enable consumers to make informed decisions when selecting vehicles based on their cybersecurity robustness.
However, many of the threat and attack models discussed in the literature do not fully address the unique cybersecurity requirements of modern automotive systems. These requirements include:
  • Interconnectivity: Modern vehicles are highly interconnected systems, comprising numerous Electronic Control Units (ECUs) that communicate with each other. Ensuring secure communication between these components is essential to prevent unauthorised access and control.
  • Safety-Critical Systems: The integration of advanced driver-assistance systems (ADAS) and autonomous driving technologies means that any cybersecurity breach could directly impact vehicle safety. Ensuring the security of these systems is paramount to protecting the safety of passengers and other road users.
  • Physical Security: Automotive systems are subject to physical threats. Attackers can potentially gain physical access to vehicles, enabling them to tamper with hardware components or inject malicious code directly.
  • Over-the-Air (OTA) Updates: Many modern vehicles support OTA updates for software and firmware. While this feature enhances convenience and functionality, it also introduces new attack vectors.
  • User Privacy: Automotive systems collect and process a significant amount of data, including personal information, driving habits, and location data. Protecting user privacy is crucial to prevent misuse of this data.
In our previous work [7], we conducted a systematic literature review of threat and attack modelling approaches pertinent to cyber-physical systems (CPSs)—systems that integrate computational and physical process—of which automotive systems are an example. Our investigation yielded several findings, including:
  • Threat modelling is commonly conducted only in early stages of system development, and does not always consider attack models.
  • Threat and attack models in the literature primarily focus on IT systems, making them challenging to apply to automotive systems.
  • Much of the existing literature does not differentiate between cybersecurity breaches and consequences in IT systems versus those in automotive systems.
  • Ambiguity persists in the literature regarding the definitions and relationships between threat modelling and attack modelling.
For instance, threat models developed using Microsoft’s STRIDE (Spoofing, Tampering, Repudiation, Information Disclosure, Denial of Service, and Elevation of Privilege) primarily focus on data flows and information security. However, STRIDE was not designed with automotive systems in mind, making its application in this context somewhat limited. Consequently, STRIDE often overlooks critical aspects of how cybersecurity breaches can directly affect automotive functions and safety. As a result, the reliability of the general outcomes from STRIDE is questionable when applied to the unique characteristics of automotive environments.
In other words, threat modelling of automotive systems often does not consider the consequences of cybersecurity attacks beyond information security. Similarly, attack models designed for traditional IT systems do not fully align with the multi-path tactics and techniques employed by attackers seeking to gain control of a vehicle. These techniques often involve a combination of physical and cyber actions, necessitating a more comprehensive and dynamic approach to security modelling.
These findings highlight the need to tailor security modeling approaches for automotive systems so that they can adequately capture the unique cybersecurity threats faced in these environments.
Contributions. In this experience report, we introduce ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling) — an integrated security modeling approach that enhances the resilience of automotive systems by dynamically updating their cybersecurity posture in response to prevailing threats, attacker tactics, and their impact on system functionality and safety. This approach addresses the existing knowledge gap in static security assessment methodologies by providing a dynamic and iterative framework. We demonstrate the effectiveness of ACTISM by applying it to an IVI system, illustrating how the model can be adapted as new threats emerge.
Unlike existing security modelling approaches, ACTISM integrates various frameworks to dynamically update security models with new information, acquired through threat intelligence and ongoing security monitoring. This keeps security models relevant against evolving attacker tactics, techniques and procedures (TTPs) even if the models were developed in early stages of system development. ACTISM also allows practitioners to apply the security models in both IT and automotive domains by using attack paths to capture the multi-path and multi-agent aspects of real-world cyber-physical attacks. In order to bolster the completeness of the security model, ACTISM integrates the following threat analysis and risk assessment frameworks:
  • Microsoft’s STRIDE methodology.
  • HEAVENS (HEAling Vulnerabilities to ENhance Software Security and Safety), a structured and model-based hazard analysis and risk assessment method specifically designed for automotive systems.
  • Attack Tree, a hierarchical diagram representing potential attack paths, illustrating how an attacker could exploit system vulnerabilities to achieve specific goals.
  • Common Vulnerability Scoring System (CVSS), an open industry standard for assessing the severity of computer system security vulnerabilities. CVSS (Common Vulnerability Scoring System): This model is also static as it provides a standardised method for evaluating the severity of vulnerabilities at a particular point in time. It does not dynamically adapt to new threats without manual intervention.
  • Threat Analysis and Risk Assessment (TARA) methodology from the ISO/SAE 21434:2021 ("Road Vehicles - Cybersecurity Engineering") Standard, jointly developed by the International Organisation for Standardisation (ISO) and the Society of Automotive Engineers (SAE), for managing cybersecurity threats and vulnerabilities throughout the life cycle of a vehicle.
Table 1 lists the differences between ACTISM and these security modelling frameworks, and their respective limitations. In contrast with other security modelling frameworks, the ACTISM framework integrates various frameworks to be dynamic, continuously adapting to new threats and updating the automotive system’s cybersecurity posture. This distinction sets it apart from the static models, providing a more robust approach to maintaining automotive cybersecurity.
Table 1: Comparison of ACTISM with other security modelling frameworks
Framework Type Limitations
ACTISM Dynamic Relies on the integration of various security modelling frameworks.
STRIDE Static Limited to specific threat categories. Does not dynamically adapt to new threats without manual intervention.
Attack Trees Static Requires manual updates and may not cover all emerging threats. Does not dynamically adapt to new threats without manual intervention.
HEAVENS Static Primarily focuses on high-level security analysis which may not cover detailed, system-specific vulnerabilities and threats. Does not dynamically adapt to new threats without manual intervention.
CVSS Static Does not dynamically adapt to new threats without manual intervention.
TARA Static Does not dynamically adapt to new threats without manual intervention.
Although frameworks such as STRIDE and CVSS were designed for IT security threats, they can also be applied to physical threats. As an example, tampering with physical components can lead to unauthorised access, data manipulation, or service disruption. As an integrative approach, ACTISM provides guidance on how both cyber and physical attacks can be modeled using these frameworks to capture the various entry points and methods attackers might use, ensuring comprehensive security coverage across both cyber and physical domains.
To demonstrate the practical application of ACTISM, we developed a comprehensive security model for an IVI system. We outline the step-by-step process of building the security model, emphasising the integration of threat intelligence and contextual information to accurately characterise and assess cybersecurity risks.
Furthermore, we demonstrate the utility of ACTISM by illustrating its adaptive response to a real-world vulnerability disclosure affecting the IVI system. This case study underscored ACTISM’s effectiveness in addressing emerging threats and evolving security challenges in automotive cybersecurity. Furthermore, the study showcased the versatility of ACTISM, demonstrating its potential applicability to other automotive systems beyond the IVI, making it a robust tool for safeguarding a wide range of vehicle technologies.
Finally, we outline avenues for future research and development in this area, including automated vulnerability management workflows.
2 Related Work
Various works have attempted to address security modelling in automotive systems.
Das et al. [4] employed the STRIDE methodology for threat modelling and utilised the SAHARA (Safety-Aware Hazard Analysis and Risk Assessment) and DREAD (Damage, Reproducibility, Exploitability, Affected Users, and Discoverability) methodologies to assess risks associated with Infotainment High Performance Computing (HPC) systems. However, their analysis did not incorporate attack modelling.
Abuabed et al. [1] proposed a framework for assessing the cybersecurity vulnerabilities of automotive vehicles. Their approach employed STRIDE to classify threats. For assessing the impact of threats, they used version 2.0 of the HEAVENS security model [9], and they evaluated attack feasibility with CVSS version 2.0. Additionally, they used a risk matrix to assess risks based on a use-case involving the Advanced Driver Assistance System (ADAS). Using a TARA approach compatible with the ISO/SAE 21434:2021 standard, the framework utilises attack trees to determine attack paths. This method facilitates a focused analysis of relevant threat vectors and potential attacks during the assessment of attack feasibility.
However, the framework has several limitations. Firstly, in the threat identification process, the authors did not include physical attack vectors, such as accessing and modifying embedded software in vehicle components, in the STRIDE Data Flow Diagram (DFD). This omission could be because of the assumption that physical attack vectors do not involve flows or exchanges of data. Notwithstanding this assumption, physical attack vectors can be exploited to facilitate data flows, such as "USB Man-in-the-Middle (USB MITM)" attacks, where an attacker intercepts and potentially alters the communication between a USB device and a host system. Consequently, physical attack vectors can be a crucial gap in the overall security design of the vehicle.
Secondly, the CVSS version 2.0 framework adopted in Abuabed et al. [1] is no longer current. The newer versions of CVSS include improvements that deliver more granularity and accuracy in the resulting risk assessment. Lastly, the framework does not adequately address threats in later stages of the life cycle. As automotive system life cycles are typically 10 to 20 years or longer, the vehicle system could become increasingly vulnerable to new types of attacks that were not anticipated during the initial security assessment.
Chlup et al. [2] proposed an automated threat modelling tool for the automotive domain. It combines threat modelling and attack tree analysis using automatically-generated attack trees and attack paths. In addition, it uses anti-patterns to perform in-depth analysis in the threat model. However, the approach does not perform attack feasibility estimation and risk assessment, which are key steps in the TARA workflow.
Other literature relevant to automotive cybersecurity includes a study by Girdhar et al. [5] who used the STRIDE methodology to create a threat model focused on post-accident cyberattack event analysis for CAV.
Additionally, Dang et al. [3] used a STRIDE framework to analyse and classify cybersecurity threats in an autonomous vehicle architecture, including the infotainment system. While the authors ranked the threats using CVSS version 3.0 scores, they did not assess the corresponding impact. The authors focused only on cybersecurity threats originating from outside the vehicle. This means that attacks that involve access to the vehicle’s internal architecture from within the vehicle are excluded in the analysis.
The current landscape of security modelling highlights several limitations. Firstly, there are differing assumptions regarding which vehicle assets fall within the trust boundaries of security models. Secondly, the feasibility and impact of cyber-physical attacks are rarely considered in security modelling. These gaps underscore the need for an iterative and cohesive framework to effectively address the evolving challenges in automotive cybersecurity.
3 Threat-informed Dynamic Security Modelling with ACTISM
We propose ACTISM - a methodology for modelling security threats and attack vectors in a dynamic, iterative manner, specifically designed for automotive systems. We will illustrate the application of ACTISM to an IVI system as a running example and apply ACTISM to a real-world case study, where we demonstrate its ability to dynamically adapt to new threats and improve the security of automotive systems. This case study provides concrete evidence of the framework’s applicability and benefits, distinguishing our research from previous work.
Current security modelling frameworks offer valuable insights but lack adaptability to evolving cyber-physical threats. ACTISM addresses this gap by introducing a framework that continuously updates the cybersecurity posture of automotive systems in response to both cyber and physical threats, attack vectors, and system vulnerabilities.
Unlike static models, ACTISM incorporates threat intelligence through continuous monitoring, which ensures that the security model remains adaptive and effective in mitigating emerging risks. This iterative approach is vital because threat and attack models created in the early stages of system life cycles can become outdated or irrelevant over time as new vulnerabilities emerge. This gap widens, particularly when cyber-physical attacks occur like those targeting keyless entry systems [15], breaching the trust boundary within security models.
ACTISM does not replace current threat and attack modelling frameworks. Instead, it streamlines their adoption and utilisation by cybersecurity practitioners. By integrating various security modelling frameworks into an iterative approach, ACTISM enhances the practicality and efficacy of addressing evolving cybersecurity threats and attacker capabilities. ACTISM comprises eight key steps as shown in Figure 1.
Figure 1: Overview of ACTISM
To enhance the practical utility of ACTISM among cybersecurity practitioners, we adopt the STRIDE methodology using a STRIDE-per-element approach, where threats are assessed against each element in the Data Flow Diagram. We choose not to use the STRIDE-per-interaction approach, as research has shown it yields a statistically significant lower level of completeness compared to the STRIDE-per-element approach [13]. Furthermore, we utilise HEAVENS version 2.0 [9] for impact assessment. To our knowledge, HEAVENS is the only framework designed to be fully compliant with the latest automotive cybersecurity standards, such as ISO/SAE 21434 and UNECE Regulation 155.
In addition, we employ CVSS version 3.1 to assess attack feasibility, which allows us to leverage contemporary cybersecurity vulnerability knowledge bases. CVSS version 3.1 provides granular metrics and scoring options, considering various factors that affect the exploitability and impact of a vulnerability. This comprehensive approach ensures that ACTISM remains robust and adaptive to the dynamic nature of cybersecurity threats in automotive systems.
4 Applying ACTISM to an In-Vehicle Infotainment (IVI) System
To illustrate the practical application of ACTISM, it is essential to understand the IVI system and its role in automotive cybersecurity.
Legacy head units offered basic features such as radio reception and CD playback, with limited connectivity. In contrast, modern IVI systems are advanced digital platforms that integrate various functionalities beyond audio and navigation. These systems serve as central hubs, providing connectivity to external networks, managing complex data streams, enabling access to vehicle diagnostics, and supporting interactive user interfaces for both drivers and passengers.
While modern IVI systems enhance convenience and operational efficiency, they also expose vehicles to increased cybersecurity risks. These systems are vulnerable to traditional threats, such as remote exploits and data breaches, as well as emerging threats like cyber-physical attacks targeting vehicle functionalities.
The complexity of modern IVI systems significantly expands their attack surface. They integrate multiple wireless communication protocols—Wi-Fi, Bluetooth, and cellular networks—and connect to external services and mobile devices, increasing potential vulnerabilities. Malicious actors can exploit weaknesses in wireless protocols or insecure network connections to gain unauthorised access. Furthermore, advanced functionalities like voice recognition and artificial intelligence add layers of complexity, creating additional attack vectors.
A dynamic approach like ACTISM allows for iterative security modelling throughout the IVI system’s life cycle, ensuring comprehensive protection against evolving threats.
4.1 Step 1: Identify assets
In the first step, identifying vehicle assets within the IVI system threat model typically involves a collaborative effort among system architects, engineers, and cybersecurity experts. This process integrates various activities, including reviewing technical specifications, examining system architecture diagrams, and analysing functional requirements documents. Additionally, considering the interconnected nature of modern IVI systems, potential dependencies and interactions between assets are also analysed to understand how vulnerabilities in one component could impact overall system security. These efforts aim to ensure a comprehensive understanding and coverage of all relevant assets within the IVI system.
In this experience report, we reference the IVI HPC system DFD from Das et al. (2024) [4] to guide our identification of IVI system assets. This initial asset identification, listed below, serves as a foundational step for conducting further detailed analysis and security modelling. It is important to note that this DFD is an example and may not encompass newer types of attacks, such as dirty road attacks that exploit the fact that autonomous vehicles rely on sensors and cameras to interpret their environment. Attackers may place objects or patterns on the road that confuse these sensors, leading to incorrect navigation decisions or even accidents.
  • In-vehicle modules
    • In-Vehicle Infotainment (IVI) System
    • Human-machine Interface (HMI)
    • Central Gateway
    • GPS Receiver
    • USB
    • Bluetooth
    • Wi-Fi
  • Data stores
    • Local maps data store
    • IVI system embedded data store
4.2 Step 2: Construct Data Flow Diagrams
In the next step, we construct and analyse DFDs to identify the vehicle trust boundary and potential threat and attack vectors. This process uses the Microsoft Threat Modelling Tool (Version 7.3.31026.3) along with the Automotive Threat Modelling Template [10] developed by the NCC Group. We conduct DFD analysis for both a typical automotive system [6] and specifically for the IVI system (Figure 2) to ensure thorough identification and analysis of potential threats and vulnerabilities.
Figure 2: Data Flow Diagram for an In-Vehicle Infotainment (IVI) System
4.3 Step 3: Model threats
Next, we use the Microsoft Threat Modelling Tool to generate threats based on the previously constructed Data Flow Diagram (DFD) for the IVI system. The tool categorises these threats according to the STRIDE threat model automatically, providing an initial classification into categories such as Spoofing, Tampering, Repudiation, Information Disclosure, Denial of Service, and Elevation of Privilege.
While the tool categorises threats by default and attempts to provide threat descriptions based on its own threat catalogues and templates, the initial output may be somewhat generic and not specifically tailored to the nuances of the IVI system’s operating environment. Therefore, our analysis involves further contextualisation and systematic review of these categorised threats.
We focus on the in-vehicle modules and data stores identified in the earlier step, ensuring that the generated threats are relevant to these components. After refining the list, we update threat descriptions to better align with the specific interfaces and nuances of the IVI system. A domain expert can perform this refinement relatively easily using the Microsoft Threat Modelling Tool, taking about 30 minutes for this example. The final curated list of threats for the IVI system is shown in Table 9 [8].
4.4 Step 4: Evaluate impact
To evaluate the impact of the identified threats, we use the HEAVENS version 2.0 framework [9] which is specifically designed for automotive risk assessments.
Our approach to impact analysis is as follows:
  • Assess impact level: By definition in the HEAVENS framework, the impact level of a specific threat is an estimate of the expected loss for various stakeholders when the threat is realised. To estimate the impact, we use four parameters directly related to the security objectives defined in HEAVENS: safety (s), financial (f), operational (o), and privacy (p).
    • Safety Impact: Refers to the safety of vehicle occupants, road users, and infrastructure.
    • Financial Impact: Encompasses all direct and indirect financial damages incurred by stakeholders.
    • Operational Impact: Pertains to operational damages that have minimal or no safety or financial consequences. For example, the loss of secondary functionalities such as entertainment systems.
    • Privacy Impact: Deals with damages resulting from privacy violations.
    Each of the four parameters is assigned one of four impact levels: None, Low, Medium, or High. For the purposes of this experience report, we used the impact level definitions from [12]. Impact levels are typically assigned by automotive domain experts who have detailed knowledge of the automotive system, its vulnerabilities, and corresponding impacts on security objectives. In our running example, we assign impact levels based on our assessment of potential safety, operational, financial, and privacy consequences. This assessment is based on our understanding from the literature, and may differ from evaluations of experts in the field.
    Next, numerical values are assigned to the impact levels. These values are chosen in a logarithmic scale (0, 1, 10, 100) to effectively capture the exponential increase in the impact levels. This helps in clearly differentiating between the levels of impact related to the security objectives, and ensures that high-impact threats are given appropriate priority.
    Table 10 [8] categorises the impact levels and parameter values according to the security objectives.
  • Calculate impact score: After estimating the individual impact levels of the security objectives and their corresponding parameter values, the overall impact score is computed using the following equation:
    $\small I=w_{s}i_{s}+w_{o}i_{o}+w_{f}i_{f}+w_{p}i_{p}$ (1)
    Normalisation is necessary to ensure that the overall impact score is on a consistent scale, allowing for fair and meaningful comparisons across different systems and contexts. This uniformity is crucial when comparing the cybersecurity posture of various systems, such as between automotive systems.
    The overall normalised impact score is computed using the following equation:
    $I_{\text{norm}}=\frac{w_{s}i_{s}+w_{o}i_{o}+w_{f}i_{f}+w_{p}i_{p}}{100*(w_{s}+w_{o}+w_{f}+w_{p})}$ (2)
    Weight values $w_{s}$, $w_{o}$, $w_{f}$, and $w_{p}$ are selected based on the specific needs of the use case being evaluated. In the case of an automotive system, the safety impact is the most severe consequence for stakeholders, such as vehicle occupants, road users, and infrastructure. Comparatively, financial, operational, and privacy impacts may have lower consequences. As such, the weights are selected as: $w_{s}=10$ and $w_{o}=w_{f}=w_{p}=1$. Equations 1 and 2 are simplified as follows:
    $I=10i_{s}+i_{o}+i_{f}+i_{p}$ (3)
    $\small I_{\text{norm}}=\frac{10i_{s}+i_{o}+i_{f}+i_{p}}{1300}$ (4)
  • Assign impact rating: Each threat is assigned an impact rating based on the impact score $I_{\text{norm}}$ of each assessed threat, matched against thresholds tabulated in Table 2, as referenced from [12].
    Table 2: Impact rating table
    Impact Score Impact Rating
    0.00 None
    $0.010.00<x<$ 0.01 Low
    $0.050.01\leq x<$ 0.05 Medium
    $0.450.05\leq x<$ 0.45 High
    $1.000.45\leq x<$ 1.00 Critical
For illustration, we show how the impact rating of Threat ID 9146 from Table 9 [8] is derived using this methodology. In this specific threat, an adversary may reverse engineer the head unit firmware to find sensitive information, resulting in information disclosure from the IVI System’s local maps data store. Using the definitions of the impact assessment table (Table 10 [8]), we evaluate $i_{s}=0$, $i_{o}=0$, and $i_{f}=0$, as there are no injuries or discernible operational and financial effects resulting from the disclosure of information. However, disclosure of personally identifiable information or passwords from the data store can result in privacy violations leading to abuses such as impersonation of the victim to perform actions with stolen identities. As such, the impact level for privacy is assessed as Medium and $i_{p}=10$. Using Equation 2, $I_{\text{norm}}\ =0.0077$, which corresponds to a Low impact rating.
Table 3: Impact rating example
Threat ID $i_{s}$ $i_{o}$ $i_{f}$ $i_{p}$ Impact Score Impact Rating
9146 0 0 0 10 0.0077 Low
4.5 Step 5: Model attacks
In this step, we use an attack tree to model attacker goals and map out attack paths associated with threat and attack vectors identified earlier using the Microsoft Threat Modelling Tool.
Since the vehicle driver is within the car trust boundary (Figure 2), we will assume the attacker personas and their goals as follows:
  • Persona A: A car thief interested in stealing sensitive data from the In-Vehicle Infotainment system.
  • Persona B: A rogue security researcher interested in manipulating vehicle functions that compromise the safety of the vehicle.
The attack tree depicted in Figure 3 illustrates the potential attack paths these two personas could employ to achieve their respective goals.
Figure 3: Attack tree for In-Vehicle Infotainment (IVI) system
4.6 Step 6: Evaluate attack feasibility
Next, we use CVSS version 3.1 to analyse and compute attack feasibility. CVSS provides a standardised method for assessing the severity of vulnerabilities based on their impact and exploitability factors, as described below.
  • Impact Metrics:
    • Confidentiality (C): Impact on the confidentiality of data, such as driver information, vehicle location, and system diagnostics.
    • Integrity (I): Impact on the integrity of automotive systems, including the accuracy and reliability of data and control signals within the vehicle.
    • Availability (A): Impact on the availability of automotive functions and services, such as the ability to start the vehicle or use navigation systems.
  • Exploitability Metrics:
    • Attack Vector (AV): How the vulnerability is accessed.
    • Attack Complexity (AC): The complexity of the attack required to exploit the vulnerability.
    • Privileges Required (PR): The level of privileges required to exploit the vulnerability.
    • User Interaction (UI): The level of user interaction required to exploit the vulnerability.
  • Impact Factor:
    • The Impact Factor depends on the scope of the vulnerability, which can be either Scope Unchanged (U) or Scope Changed (S). If the vulnerability affects other system components, the Scope is considered Changed (S), otherwise, it is Unchanged (U).
In this experience report, instead of using the CVSS base score, which is common in the literature, we use the CVSS temporal score to incorporate and emphasise the temporal aspects of the vulnerability, such as exploit code maturity, remediation level, and report confidence. These are shown at Equations 5 and 6.
CVSS Base Score=Round up(0.6×Impact+0.4×Exploitability1.5Impact $\text{CVSS Base Score}=\text{Round up}\left(\frac{{0.6\times\text{Impact}+0.4\times\text{Exploitability}-1.5}}{{\text{Impact Factor}+0.8}}\right)$ (5)
CVSS Temporal Score=Round up(CVSS Base Score×Exploit Code Maturity×Remediation Level×Report $\begin{split}\text{CVSS Temporal Score}=\text{Round up}(\text{CVSS Base Score}\times\text{Exploit Code Maturity}\times\\ \text{Remediation Level}\times\text{Report Confidence})\end{split}$ (6)
Next, the CVSS temporal scores for the identified threats are computed using the CVSS version 3.1 calculator (https://nvd.nist.gov/vuln-metrics/cvss/v3-calculator), then mapped to qualitative severity ratings using Table 4 as referenced from the CVSS 3.1 Specification Document. The severity ratings, which we refer to as attack feasibility in this experience report, are tabulated in Table 5.
Table 4: Qualitative severity rating scale
CVSS Score Severity Rating (Attack Feasibility)
0.0 - 3.9 Low
4.0 - 6.9 Medium
7.0 - 8.9 High
9.0 - 10.0 Critical
Table 5: Attack feasibility based on CVSS version 3.1
Threat ID AV AC PR UI S C I A E RL RC CVSS Score Attack Feasibility
9158 A L N N C N L H P U U 7.1 High
9132 A H N N C N H H U U U 6.7 Medium
9089 P L N R U H H H U U U 5.6 Medium
9144, 9060 A L L R U H H H U U U 6.2 Medium
9146 P H N R U H L N U U U 3.9 Low
4.7 Step 7: Evaluate risks
In this step, we evaluate the cybersecurity risks of the IVI system using a 5x5 risk matrix illustrated in Figure 4. In the context of automotive impact assessment, where the stakes can be high and risks multifaceted, a 5x5 risk matrix provides the depth and granularity needed to accurately assess and manage risks, compared to a simpler 3x3 risk matrix. The risk ratings for the identified threats are tabulated in Table 6.
Figure 4: 5x5 risk matrix
Table 6: Risk ratings based on attack feasibility and impact ratings
Threat ID Description Attack Method Source Attack Feasibility Impact Rating Risk Rating
9132 Cause a denial-of-service of the central gateway, affecting vehicle functions Inject or flood invalid data to the central gateway causing it to crash or stop. IVI System High Critical Extreme
9159 Cause gateway to misbehave, crash or stop, affecting vehicle functions Exploit vulnerabilities, through outdated software or configuration weaknesses. IVI System Medium High High
9089 Elevate privileges on the IVI System (via USB) Exploit vulnerabilities, through outdated software or configuration weaknesses, via USB removable media. USB Medium Low Medium
9144, 9060 Elevate privileges on the IVI System (via Bluetooth or Wi-Fi) Exploit vulnerabilities, through outdated software or configuration weaknesses, via Bluetooth or Wi-Fi. Bluetooth or Wi-Fi Medium Low Medium
9146 Access sensitive data e.g. passwords from IVI System Embedded Data Store Reverse engineer the head unit firmware to find keys to access the IVI System Embedded Data Store. Physical Access Low Low Medium
These risk ratings serve as a guideline for the Original Equipment Manufacturer (OEM) of the vehicle or the manufacturer of the IVI system. They can use these ratings to prioritise the implementation of mitigation controls throughout the system life cycle. This approach ensures that adequate attention and resources are allocated to address the most critical cybersecurity concerns identified through this comprehensive analysis.
4.8 Step 8: Review, update, and iterate
In this final step, we continuously review the cybersecurity analysis and update the risk assessment based on new threat information, insights or changes in the automotive system environment. The updated risk assessment helps refine existing mitigation strategies and enhances security controls against emerging cybersecurity threats. This iterative cycle of review and improvement is crucial for maintaining the automotive system’s long-term cybersecurity posture.
5 Updating the Security Model After a Vulnerability Disclosure
ACTISM enhances security modelling by introducing a dynamic and iterative methodology, resulting in a threat-informed, evolving security model. This approach addresses evolving cybersecurity threats that necessitate the incorporation of new knowledge gained from security monitoring and threat intelligence throughout the system’s life cycle.
We now apply this methodology to a real-world example using the Tesla Electric Vehicle (EV) jail-break vulnerability disclosure [14]. Although this vulnerability is specific to the Tesla EV platform, ACTISM is designed to be applicable to all automotive systems (including non-EVs) and their relevant vulnerabilities.
After gathering technical details from the vulnerability disclosure, we review the IVI system security model developed in the earlier section and update the CVSS metrics as necessary.
For this vulnerability disclosure, there is no need to include additional assets in the security model or to revise the DFD (Figure 2) and attack tree (Figure 3). Although there is an opportunity to enhance the DFD and attack tree to include new attack vectors, such as cyber-physical attacks like keyless entry systems [15], we will not do so in this example. However, it is critical to reassess the CVSS metrics associated with Threat ID 9146, which correlates with the vulnerability disclosure.
The CVSS metrics corresponding to Threat ID 9146 are updated as follows:
  • Scope (S) From Unchanged (UC) to Changed (C), as successful exploitation of the vulnerability could grant an attacker full control over the IVI system, potentially enabling them to access and manipulate certain vehicle functions.
  • Integrity (I) From Low (L) to High (H), as successful exploitation of the vulnerability could grant an attacker root-level read and write access to the IVI file system.
  • Availability (A) From None (N) to High (H), as an attacker could deny access to resources in the IVI system.
  • Exploit Code Maturity (E) From Unproven (U) to Proof-of-Concept (P), in view of the disclosure from the security researchers.
  • Remediation Level (RL) From Unavailable (U) to Temporary Fix (T), as the AMD Trusted Platform Module (TPM) in affected IVI systems could still be susceptible to voltage glitching attacks.
  • Report Confidence (RC) From Unknown (U) to Confirmed (C), in view of the disclosure from the security researchers.
With these updated parameters, the attack feasibility increased from 2 (Low) to 3 (Medium), as shown in Table 7.
Table 7: Updated attack feasibility following vulnerability disclosure
Threat ID AV AC PR UI S C I A E RL RC CVSS Score Attack Feasibility
9158 A L N N C N L H P U U 7.1 High
9132 A H N N C N H H U U U 6.7 Medium
9089 P L N R U H H H U U U 5.6 Medium
9144, 9060 A L L R U H H H U U U 6.2 Medium
9146 P H N R C H H H P T C 6.4 Medium
Besides attack feasibility, the impact assessment of Threat ID 9146 in the security model also needs updating. Using the definitions from the impact assessment table (Table 10 [8]), we estimate the impact on safety to increase from None to Low ($i_{s}=10$), as there is now a possibility of light and moderate injuries, should the attacker access and manipulate certain vehicle functions via the vulnerability. We also expect the operational impact to increase from None to Low ($i_{o}=10$), as there may be a degradation of primary vehicle functions. The financial impact is also expected to increase from None to Low ($i_{f}=10$), as the financial damage from injury claims may be substantial but not threatening the existence of the manufacturer. Lastly, the privacy impact is expected to increase from Low to High ($i_{p}=100$), as the privacy violation of multiple vehicle owners and drivers may lead to extensive media coverage and severe loss of market share. Using Equation 2 to compute the updated normalised impact score, $I_{\text{norm}}\ =0.1692$, which corresponds to a High impact rating (Table 8).
Table 8: Updated impact rating following vulnerability disclosure
Threat ID $i_{s}$ $i_{o}$ $i_{f}$ $i_{p}$ Updated Impact Score Updated Impact Rating
9146 10 10 10 100 0.1692 High
With the updated attack feasibility (Medium) and impact rating (High), the overall risk rating for Threat ID 9146 increased from Medium to High, in accordance with the 5x5 risk matrix (Figure 4). This ability to perform timely updates to the risk assessment is crucial for maintaining robust cybersecurity in automotive systems. Ultimately, it empowers stakeholders to adopt additional precautions and improve the cybersecurity resilience of their vehicles throughout their life cycle.
6 Future directions
While ACTISM provides an approach to dynamically updating security models with both cyber and physical threats, it is not without its limitations. One of the primary challenges is how the accuracy of the model can be validated given the constantly evolving nature of threats and the potential for incomplete or inaccurate threat intelligence data. Moreover, while ACTISM’s adaptability to new threats is a strength, it can also pose a challenge in maintaining the models’ integrity and consistency over time. Despite these limitations, the potential future directions outlined below can address these challenges and enhance the robustness of ACTISM.
Ongoing threat intelligence gathering, including real-time monitoring and tracking of vulnerability disclosures, is crucial for keeping the security model current. Establishing a comprehensive repository of vulnerability attributes, stored in a standardised format such as XML or JSON, can enable automation and streamline the updating process. Additionally, formalising security models using Domain-Specific Languages (DSLs) like Automation Markup Language (AutomationML) and Meta Attack Language (MAL) can provide clearer descriptions of threats and vulnerabilities.
Incorporating adversary tactics from the MITRE ATT&CK knowledge base can further enhance these models by offering insights into real-world attack strategies. Moreover, leveraging machine learning approaches to prioritise vulnerabilities based on their impact, utilising historical data and threat intelligence, can significantly improve automotive cybersecurity.
These potential future directions can ensure that automotive cybersecurity remains resilient and adaptive to emerging cybersecurity challenges.
7 Conclusion
The cybersecurity landscape for automotive systems is rapidly evolving, with attackers increasingly targeting vulnerabilities that threaten data integrity, functionality, and safety. Traditional static security assessment methodologies, such as STRIDE, often fall short in addressing the dynamic and multi-faceted nature of these threats. There is a clear knowledge gap in the ability to model and respond to security threats in a dynamic, iterative manner tailored specifically for automotive systems.
In response, we introduced ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling), an integrated approach designed to enhance the resilience of automotive systems by dynamically updating their cybersecurity posture. ACTISM effectively responds to evolving cyber-physical threats and attacker tactics, ensuring that security models remain robust and relevant throughout a system’s lifecycle.
To illustrate the practical application of ACTISM, we developed a comprehensive security model for an IVI system. This process included identifying key assets, conducting thorough threat modelling, and performing risk assessments, ultimately resulting in a robust security framework. A real-world case study involving Tesla’s IVI system further demonstrated ACTISM’s ability to adapt to emerging threats and underscored its potential for broader applicability across various automotive technologies.
In conclusion, ACTISM addresses the critical knowledge gap in automotive cybersecurity by providing a proactive, adaptable approach to security modelling. This approach not only tackles current challenges but also prepares for the future landscape of automotive technology.